Determine what is asked. Show all pertinent parts of the solution (GUFSA), make sure to express your answer with correct number of significant figures.


  1. Velocity at t = 6s, considering that the initial velocity is 4m/s
1
Expert's answer
2021-10-04T13:34:14-0400

1) Put that the acceleration is 1m/s².

The velocity after time t is

"v=v_0+at,\\
v=4+1\cdot6=10\text{ m/s}."